OpenStudy (anonymous):

Solve the equation. = 5 A. x = -4 B. x = -12 C. x = -20 D. x = -28

OpenStudy (anonymous):

@mathshelp14

OpenStudy (anonymous):

plz help @e.mccormick

OpenStudy (e.mccormick):

Something is missing.

OpenStudy (anonymous):

oh yeah just a minute

OpenStudy (anonymous):

. PREALG4M 3.3.1-15+ Solve the equation. = 5 A. x = -4 B. x = -12 C. x = -20 D. x = -28

OpenStudy (anonymous):

is that what was missing?

OpenStudy (e.mccormick):

Nope. =5 is not an equation. It is half of one.

OpenStudy (anonymous):

no no n o n just a minute

OpenStudy (anonymous):

PREALG4M 3.3.1-15+ Solve the equation. = 5

OpenStudy (anonymous):

it wont show the other side

OpenStudy (e.mccormick):

Well, type it.

OpenStudy (anonymous):

x/-4

OpenStudy (anonymous):

but its supposed to go up and down

OpenStudy (e.mccormick):

\(\dfrac{x}{-4}=5\)

OpenStudy (anonymous):

yeah

OpenStudy (e.mccormick):

OK. So, what can you multiply both sides by to get the -4 over to the right of the =?

jaynator495 (jaynator495):

the answer is x=-20 because scince there both negative... its like deviding 4 by 20 which is 5 so -4/-20=5 i hope this helped ;), -Jaynator495

OpenStudy (anonymous):

um

OpenStudy (anonymous):

thank you

jaynator495 (jaynator495):

your welcome ;)

OpenStudy (e.mccormick):

This is the process I was trying to get you to see: \(\dfrac{x}{-4}=5\) \(\dfrac{-4}{1}\cdot\dfrac{x}{-4}=5\cdot\dfrac{-4}{1}\) \(\dfrac{\cancel{-4}}{1}\cdot\dfrac{x}{\cancel{-4}}=\dfrac{(5)\cdot(-4)}{1}\) \(x=\dfrac{-20}{1}\) \(x=-20\) The key is that multiplication and division undo each other. So when you have a division question, you can use multiplication to both sides to undo it and find the answer or a step towards the answer.